Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $1.25 coming 3 years from today. The dividend should grow rapidly - at a rate of 15% per year - during Years 4 and 5, but after Year 5, growth should be a constant 5% per year. If the required return on Computech is 15%, what is the value of the stock today? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Dividend for 3rd year (D3)=   1.25
Growth rate for 4th and 5th years=   15.00%


Dividend for 4th year (D4)= d3*(1+g)  
1.25*(1+15%)=   1.4375


Dividend for 5th year (D5)= 1.4375*(1+15%)=   1.653125
Constant growth rate thereafter=   5%


D6= 1.653125*(1+5%)=    1.735781250
Required rate of return =   15%


Terminal value at end of Year 5= D6/(required Return - growth rate)  
1.73578125/(15%-5%)  
17.3578125  
  
Current price = Present Value of 3-5 year dividends + Present Value of terminal value  
(D3/(1+ke)^3) + (D4/(1+ke)^4)+(D5/(1+ke)^5 )+(Terminal value/(1+ke)^5)  
=(1.25/(1+15%)^3)+(1.4375/(1+15%)^4)+(1.653125/(1+15%)^5)+(17.3578125/(1+15%)^5)  
=$11.09558642  
So, Current price of stock is    $11.10

Please thumbs up